OpenStudy (anonymous):

Simplify the expression: [(14 + 2) • 4 – 1] • 6 114 186 288 378

OpenStudy (anonymous):

What do you think the answer is?

OpenStudy (anonymous):

well multiplication goes first

OpenStudy (anonymous):

but parenthesis does to

OpenStudy (anonymous):

You solve the parenthesis first.

OpenStudy (anonymous):

So you have 16 * 4 - 1

OpenStudy (anonymous):

D is my aswer

OpenStudy (anonymous):

No, its actually C.

OpenStudy (anonymous):

how is tht possible 16 x 4=64 - 1 = 63x 6=378

OpenStudy (anonymous):

[(14 + 2) • 4 – 1] • 6 14 + 2 = 16 4 - 1 = 3 16 * 3 = 48 48 * 6 = 288

OpenStudy (anonymous):

youre suppose to do everything in parenthesis first

OpenStudy (anonymous):

You follow PEMDAS. Parenthesis. Exponents, Multiplication, Division, Addition, Subtraction.

OpenStudy (anonymous):

ya parenthesis first

OpenStudy (anonymous):

u even did 16*4-1 which is 63

OpenStudy (anonymous):

times 6, 378

jigglypuff314 (jigglypuff314):

D is correct :)

OpenStudy (anonymous):

Hmmm your right. I totally did that wrong. Sorry bout that :)

OpenStudy (anonymous):

im so glad i stuck to my answer

OpenStudy (anonymous):

Lol so am I :)

OpenStudy (anonymous):

thnx for the help though

OpenStudy (anonymous):

Your welcome. I'm gonna start double checking myself before I answer :P